# Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning

### v2

@@ Abstract
Drawing inspiration from human learning behaviors, this work proposes a novel approach to mitigate catastrophic forgetting in Prompt-based Continual Learning models by exploiting the relationships between continuously emerging class data. We find that applying human habits of organizing and connecting information can serve as an efficient strategy when training deep learning models. Specifically, by building a hierarchical tree structure based on the expanding set of labels, we gain fresh insights into the data, identifying groups of similar classes could easily cause confusion. Additionally, we delve deeper into the hidden connections between classes by exploring the original pretrained model’s behavior through an optimal transport-based approach. From these insights, we propose a novel regularization loss function that encourages models to focus more on challenging knowledge areas, thereby enhancing overall performance. Experimentally, our method demonstrated significant superiority over the most robust state-of-the-art models on various benchmarks.
@@ 1 Introduction
Continual Learning (CL) [31, 16] is a research direction focused on realizing the human dream of creating truly intelligent systems, where machines can learn on the fly, accumulate knowledge, and operate in constantly changing environments as a human’s companion. Despite the impressive capabilities of A.I systems, continual learning remains a challenging scenario due to the tendency to forget obtained knowledge when facing new ones, known as catastrophic forgetting [5]. In dealing with this challenge, traditional CL methods often rely on storing past data for replaying during new tasks, which can raise concerns about memory usage and privacy. To overcome this limitation, recent methods proposed leveraging the generalizability of pre-trained models [7, 10] as frozen backbones to solve sequences of CL tasks [34, 25, 15].
@@
While these pre-trained-based methods have demonstrably achieved impressive results, only consider forgetting caused by changes in learned prompts or differences between the (prompt-based) models chosen for training and testing [30, 27, 37]. Further completing those arguments, we show that forgetting of old knowledge also comes from the uncontrolled growth of new classes in the latent space. That is, models are confused in distinguishing between old and new classes, which many methods overlook when training tasks independently [25, 35]. Furthermore, we find that current approaches only utilize limited information from the training dataset and treat class labels equally during training, resulting in missing opportunities to further enhance model representations and mitigate forgetting more effectively.
@@
In addition, we find that human natural learning behavior has many valuable aspects, especially the habit of analyzing data, organizing them in a meaningful way, and finding connections between old and new knowledge [24, 2, 26, 19], thereby improving the ability to understand, remember, and reproduce information. Inspired by this, we investigate the characteristics of current common benchmark datasets as well as the behavior of pre-trained-based CL models, showing that the incoming data classes over time can always be categorized into consistent groups. Each such group usually includes class data with similar semantic characteristics, which is easy to get confused by models and should be paid special attention to during training.
Therefore, we propose a training strategy that constantly considers emerging data in groups, following a hierarchical tree-like taxonomy developed based on expert/domain knowledge. During training new tasks, the model references information from old classes in the tree. Especially, the feature extractor is encouraged to maximally contrast and distinguish concepts/labels within the same group, also promoting the learning of common features that can be transferred to new concepts/labels in the same group in future tasks. This strategy not only mitigates forgetting when new classes emerge but also consolidates domain-specific knowledge. Besides, by considering human learning habits, we also find that individuals with better foundations will absorb knowledge faster and more effectively. Thus, we propose an optimal-transport-based technique to further exploit prior knowledge from pretrained models, where the behavior of this starting model will reveal a new view of the relationships between classes. Moreover, we propose a testing strategy where the characteristics of each group aid in identifying data samples, especially those located at the boundaries between different leaf groups of the taxonomy.
@@
Contribution. We name our method as Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning (TCL), and summarize our main contributions as follows:
@@
  • We introduce a new perspective to explain the reason for catastrophic forgetting in pretrained-based Continual learning models, which potentially comes from the uncontrolled growth of incoming classes on the latent space.
@@
  • Originating from the research findings of Cognitive Science, we propose a novel approach to reduce forgetting by exploiting relationships between data. By dynamically building label-based hierarchical taxonomies and leveraging the initial behavior of the pretrained model, we can identify the challenging knowledge areas that further focus, and determine how to learn them during the sequence of tasks. Based on this taxonomical structure, our testing strategy further improves model performance.
@@
  • We empirically evaluate the effectiveness of our method against current state-of-the-art pre-trained-based baselines across various benchmarks.
@@
Organization. The rest of the paper is structured as follows. In Section 2, we present related work. Then in Section 3, we formulate the problem and introduce a new perspective to explain the cause of forgetting in CL models. Section 4 transitions from the motivation provided by Cognitive Science insights to the proposed training and testing strategy, emphasizing the importance of exploiting relationships between class data. Section 5 presents the experimental results, and finally, we discuss the limitations and suggest future directions in Section 6.
@@ 2 Related work
This line of work exploits the power of pre-trained backbone to quickly adapt to the sequence of downstream tasks by updating just a small number of parameters (prompts). Initial work like [35, 34, 25] typically assign a set of prompts to tasks, enhancing the adaptability of the backbone to downstream tasks. However, the absence of explicit constraints can lead to feature overlapping between classes from different tasks. Therefore, recent methods employ some types of contrastive loss [30, 14] or utilize Vision Language models [32, 21] to better separate features from tasks. However, they treat all classes equally during training, missing the opportunity to learn in challenging areas where classes have many similarities and are easily confused. In this work, we propose a novel approach to exploit the relationships within data, allowing the model to recognize groups of these classes, and develop a deeper understanding of the respective knowledge areas, thereby reducing forgetting and enhancing its ability to learn new tasks.
3 Class Incremental Learning and forgetting in prompt-based models
3.1 Problem formulation and Notations
@@
We consider the Class Incremental Learning setting [38, 17, 30], where a model has to learn from a sequence of $T$ visual classification tasks without revisiting old task data during training or accessing task IDs during inference. Each task $t\in\{1,...,T\}$ has a respective dataset $\mathcal{D}_{t}$, containing $n_{t}$ i.i.d. samples ${(\bm{x}_{t}^{i},y_{t}^{i})}_{i=1}^{n_{t}}$. Let $D_{c}=(X_{c},Y_{c})$ denote the data corresponding to the class label $c$.
@@
In this work, we design our model as a composition of two components: a pre-trained ViT backbone $f_{\Phi}$ and a classification head $h_{\psi}$. That is, we have the model parameters $\theta=(\Phi,\psi)$. Similar to other existing prompt-based methods, we incorporate into the pre-trained ViT a set of prompts $\bm{P}$. We denote the overall network after incorporating the prompts as $f_{\Phi,\bm{P}}$.
3.2 Forgetting in Prompt-based Continual Learning
In CL models, changes in the dataset, including inputs $\bm{X}$ in the input space $\mathcal{X}$ and labels $\bm{Y}$ in the label space, $\mathcal{Y}$ lead to the changes of model’s behavior (feature shift) and thus model’s performance on previously learned tasks to decrease significantly (i.e., catastrophic forgetting). Current prompt-based CL methods, which leverage the power of pretrained models, attribute forgetting/feature shift either to (I) changes in parameters from the backbone when using the common prompt pool $\bm{P}$ for all tasks [35, 34] or to (II) the inherent mismatch between the models used at training and testing. Specifically, let $\hat{t}(\bm{x})$ and $t(\bm{x})$ as the chosen promptID and the ground-truth promptID for $\bm{x}$, respectively. We may have $f_{\hat{\theta}_{t}}=f_{\Phi,\bm{P}_{\hat{t}}}\neq f_{{\theta}_{t}}=f_{\Phi,\bm{P}_{{t}}}$ because there are chances that the promptID $\hat{t}\neq t$, where $\hat{t}$ is predicted by pretrained backbone $f_{\Phi}$ (Figure 1(a)), as discussed and analyzed in [37, 27]. To complement these views, below we provide an empirical study to offer new insights about the reason for forgetting in this type of model, which arises from overlapping between old and new class representations.
Firstly, to completely eliminate concerns about changing learned parameters, we consider methods that propose using a distinct set of prompts $\bm{P}_{t}$ to a specific task $t$. Then, the remaining potential factor of forgetting by feature shift is the difference between the prompt chosen at inference time and the one used during training (i.e., $\bm{P}_{t}\neq\bm{P}_{\hat{t}}$). Thus, we conduct experiments on HiDE [30] (i.e., the latest SOTA in prompt-based CL) to measure the differences between the features formed when using these two prompts. Particularly, we consider $\bm{x}\in D_{1}$ that belongs to the first task and measure the L2 Wasserstein distance $W_{2}(Q,\hat{Q})$ [11] between $Q$ (i.e., the latent distribution corresponding to $\bm{P}_{t}$, which consists of $f_{\Phi,\bm{P}_{{t}(\bm{x})}}(\bm{x})$ for $\bm{x}\sim D_{1}$) and $\hat{Q}$ (i.e., the latent distribution corresponding $\bm{P}_{\hat{t}}$ after learning the last task, consists of $f_{\Phi,\bm{P}_{\hat{t}(\bm{x})}}(\bm{x})$ for for $\bm{x}\sim D_{1}$). The results in Table 2 show that the difference of these distributions is apparently negligible in many cases when using pretrained backbone Sup-21K.
(a) PromptID accuracy
Sup21K
Split-Imagenet-R $3.23\times 10^{-6}$
Split-CIFAR100 $2.46\times 10^{-6}$
Split-CUB-200 $2.15\times 10^{-6}$
Table 1: Distribution shift.
(a) Within task accuracy on $D_{1}$
(b) True accuracy on $D_{1}$
(c) Overall accuracy
[…]
@@ 3.2 Forgetting in Prompt-based Continual Learning
Figure 2: Empirical study about forgetting (HiDE). (a) Average accuracy of promptID prediction for all tasks; (b) Accuracy of the first task over time, using classification head $s_{1}$ (Sup-21K); (c) Accuracy of the first task over time, using classification head $h_{\psi}$ (Sup-21K); (d) Average accuracy on all tasks so far, after learning each task (Sup-21K). Table 2 (Distribution shift) reports L2 Wasserstein distance between the latent distributions corresponding to $\bm{P}_{t}$ and $\bm{P}_{\hat{t}}$ of data task 1, after learning the final task.
For a closer look, besides the main classification head $h_{\psi}$ used for all classes so far, at the end of each task $t$, we set up a specific classifier $s_{t}$ optimized on the frozen latent space of $D_{t}$ and then kept fixed. From now on, we refer to the accuracy measured on $D_{t}$ using $s_{t}$ as ’within task accuracy’, and the accuracy using $h_{\psi}$ as ’true accuracy’ of this task. The results in Figure 2(a), show that within task accuracy of the first task stays almost unchanged, which concurs with our observation on the negligible shift between $f_{\Phi,\bm{P}_{\hat{t}(\bm{x})}}(\bm{x})$ and $f_{\Phi,\bm{P}_{t(\bm{x})}}(\bm{x})$. Meanwhile, we observe a significant decrease in the corresponding true performance in Figure 2(b), raising the question of whether we have overlooked additional factors contributing to final forgetfulness (Figure 2(c)), beyond the issue of selecting incorrect task prompts during inference.
@@
Considering the inference feature space of $f_{\Phi,\bm{P}_{\hat{t}(\bm{x})}}(\bm{x})$, we can see that as more tasks arrive, the number of classes increase, making the space fuller and increasing the possibility of overlap between class distributions. To demonstrate this point, we provide t-SNE visualization of class representations in Figure 4 and the respective illustration in Figure 5. In particular, after Task 1, we have representations of "oak tree", "mouse" and "porcupine" located in quite separate locations. However, when Task 2 and then Task 3 arrive, the appearance of "willow tree" and "pine tree" makes the latent space become fuller, and "oak tree" no longer maintains the separation from the remaining classes as before and even its representation may even be misassigned to other classes, leading to a remarkable drop in performance.
@@
Therefore, another key cause of catastrophic forgetting in prompt-based continual learning (CL) that should be recognized is the addition of new classes, which gradually fills the latent space and overlaps with existing ones. This overlap causes confusion in distinguishing between classes, thus reducing performance over time. While existing CL methods emphasize the importance of representation learning to keep classes distinct, none explicitly acknowledge the overlap between new and old tasks as a source of forgetting. Recognizing this motivates us to propose a novel method, focusing on identifying easily confused class pairs, thereby reducing forgetting and improving performance.
@@ 4 Proposed method
In the previous section, we noted that increased overlap in data representations as more tasks arrive is one of the main reasons for greater confusion in predictions, leading to performance degradation. It is crucial to identify easily confused classes/concepts to effectively enhance their distinguishability. Inspired by cognitive science studies [24, 2, 19], showing that organizing concepts in a tree-like taxonomy of visually and semantically related items aids memory and retrieval, we propose using expert/domain knowledge to structure the concepts/labels of continual learning (CL) tasks in a hierarchical taxonomy. Interestingly, we find that concepts/labels within the same group in this taxonomy tend to be visually and semantically similar, potentially causing more overlap in the latent space and confusion for the CL classifier. Motivated by this observation, we propose group-based contrastive learning to maximize the distinguishability of these concepts/labels.

### v3

@@ Abstract
Humans perceive the world as a series of sequential events, which can be hierarchically organized with different levels of abstraction based on conceptual knowledge. Drawing inspiration from this human cognition aspect, we explore how human tendencies to organize and connect information can enhance the training of deep learning models. Specifically, we propose a novel approach to mitigate catastrophic forgetting by exploiting the relationships between continuously emerging class data. By building a hierarchical tree structure based on the expanding set of labels, we gain fresh insights into the data, identifying groups of similar classes could easily cause confusion. Additionally, we delve deeper into the hidden connections between classes by exploring the original pretrained model behavior through an optimal transport-based approach. From these insights, we propose a novel regularization loss function that encourages models to focus more on challenging knowledge areas, thereby enhancing overall performance. Experimentally, our method demonstrated significant superiority over state-of-the-art methods on various continual learning benchmarks. Our code is available at https://anonymous.4open.science/r/HierC-089B/.
*Hoang Phan and Minh Le contributed equally to this work
@@ 1 Introduction
Continual Learning (CL) [45, 27] is a research direction that focuses on realizing the human dream of creating truly intelligent systems, where machines can learn on the fly, accumulate knowledge, and operate in constantly changing environments as a human’s companion. Despite the impressive capabilities of A.I systems, continual learning remains a challenging scenario due to the tendency to forget obtained knowledge when facing new ones, known as catastrophic forgetting [12]. In dealing with this challenge, traditional CL methods often rely on storing past data for replaying during new tasks [28, 6], which can raise concerns about memory usage and privacy. Besides, prior work shows that replay methods result in overfitting and poor generalization [28, 42, 7]. To overcome these limitations, recent methods proposed leveraging the strong generalization ability of pretrained models [14, 19] to solve sequences of CL tasks. A notable line of work is the prompt-based approach [48, 38, 25], where a small set of learnable prompts is injected to pretrained backbones for adapting emerging tasks over time.
@@
While these prompt-based methods have demonstrably achieved impressive results, they only consider forgetting caused by changes of common parameters across tasks during learning [49, 48] or the inherent mismatch between the chosen prompts at training and testing [44, 40, 53, 38]. In this work, we further complement these views by showing that the forgetting of old tasks can potentially come from the uncontrolled overlap between old and new emerging class representations in the latent space. That is, models become more confused in distinguishing these classes, resulting in performance degradation w.r.t previous tasks over time (i.e., forgetting). Furthermore, we find that existing methods only utilize limited information from the training dataset and often treat classes equally during training. Consequently, they overlook the opportunities to enhance the distinguishability of models, especially between the class representations of old and new tasks, and thus, hinder the models’ ability to mitigate forgetting [11, 21].
@@
In addition, we discover that human natural learning behavior has many valuable aspects, especially the habit of analyzing data, organizing them in a meaningful way, and finding connections between old and new knowledge [37, 3, 39, 30]. Inspired by these practices, we investigate the characteristics of current common benchmark datasets, as well as the behavior of pretrained models. Our findings reveal that the data can always be categorized into consistent groups, regardless of their arrival times. Each group usually includes classes with similar semantic characteristics that models may confuse and, thus, should be paid special attention to during training. Identifying these kinds of groups will benefit in improving model representation learning, as well as aiding in addressing catastrophic forgetting.
Therefore, we propose a training strategy that constantly considers emerging data in groups, following a hierarchical tree-like taxonomy developed based on expert knowledge. In particular, during the training on new tasks, the feature extractor is trained to distinguish all classes so far, especially focusing on classes within the same group. We observe that images belonging to concepts/labels within the same leaf group in the taxonomy share strong visual and semantic correlations, leading to overlapping representations in the latent space, which compromises performance. By encouraging the feature extractor to separate and contrast their representations more distinctly, we can reduce the overlap of easily confused classes, thereby further improving performance. This strategy not only mitigates forgetting when new classes emerge but also consolidates domain-specific knowledge in each of these leaf groups.
@@
Contribution. We introduce a method named Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning (TCL). Our main contributions are summarized as follows:
@@
  • We conduct empirical analysis to suggest a new cause of catastrophic forgetting in prompt-based continual learning models, which inherently comes from the uncontrolled growth of incoming classes on the latent space.
@@
  • Originating from the research findings of cognitive science, we propose a novel approach to reduce forgetting by exploiting relationships between data. By dynamically constructing label-based hierarchical taxonomies and leveraging the prior knowledge of pretrained models via an optimal transport-based approach, we can identify the challenging knowledge areas that need further focus during the sequence of tasks.
@@
  • We empirically evaluate the effectiveness of our method against current state-of-the-art pretrained-based baselines across various continual learning benchmarks.
@@
Organization. The rest of the paper is structured as follows. In Section 2, we present related work. Then in Section 3, we formulate the problem and introduce a new perspective to explain the cause of forgetting in prompt-based CL models. Section 4 transitions from the motivation provided by Cognitive Science insights to the proposed training strategy, emphasizing the importance of leveraging relationships between class data. Section 5 presents the experimental results, and finally, we discuss the limitations and suggest future directions in Section 6.
@@ 2 Related Work
This line of work exploits the power of pretrained backbone to quickly adapt to the sequence of downstream tasks by updating just a small number of parameters or prompts. Initial work like [49, 48, 38] typically assign a set of prompts to tasks, enhancing the adaptability of the backbone to downstream tasks. However, the absence of explicit constraints can lead to feature overlapping between classes from different tasks. Therefore, recent methods employ some types of contrastive loss [44, 24] or utilize Vision Language models [46, 33] to better separate features from tasks. However, they treat all classes equally during training, missing the opportunity to learn in challenging areas where classes have many similarities and are easily confused. In this work, we propose a novel approach to exploit the relationships within data, allowing the model to recognize groups of these classes, and develop a deeper understanding of the respective knowledge areas, thereby reducing forgetting and enhancing its ability to learn new tasks.
3 Preliminary Analysis
In this section, we first present the problem formulation of prompt-based continual learning in Section 3.1. Following this, we analyze the sources of forgetting in these methods in Section 3.2.
3.1 Problem formulation
@@
We consider the Class Incremental Learning setting [54, 28, 44], where a model has to learn from a sequence of $T$ visual classification tasks without revisiting old task data during training or accessing task IDs during inference. Each task $t\in\{1,...,T\}$ has a respective dataset $\mathcal{D}_{t}$, containing $n_{t}$ i.i.d. samples ${(\bm{x}_{t}^{i},y_{t}^{i})}_{i=1}^{n_{t}}$.
@@
In this work, we design our model as a composition of two components: a pretrained Vision Transformer (ViT) [9] backbone $f_{\Phi}$ and a classification head $h_{\psi}$. That is, we have the model parameters $\theta=(\Phi,\psi)$. Similar to other existing prompt-based methods [44, 38], we incorporate into the pretrained ViT backbone a set of prompts $\bm{P}$. We denote the overall network after incorporating the prompts as $f_{\Phi,\bm{P}}$.
Dataset Sup-21K DINO-1K
Split-Imagenet-R 3.82 80.31
Split-CIFAR100 1.32 48.43
Split-CUB-200 2.04 63.56
(a) Distribution shift of representations corresponding to $\mathcal{D}_{1}$ after learning sequence of tasks.
(b) Within task accuracy on $\mathcal{D}_{1}$ (Sup-21K)
(c) True accuracy on $\mathcal{D}_{1}$ (Sup-21K)
@@
Figure 1: Empirical study about forgetting of HiDe-Prompt. Table (a) reports L2 Wasserstein distance between the latent distributions corresponding to $\bm{P}_{t}$ and $\bm{P}_{\hat{t}}$ of data task 1, after learning the final task; Figure (b) Accuracy of the first task over time, using classification head $s(\bm{x})$; Figure (c) Accuracy of the first task over time, using classification head $h_{\psi}$.
3.2 Forgetting in Prompt-based Methods
In the context of Continual Learning, forgetting refers to the phenomenon in which performance on previously learned tasks decreases over time. Current prompt-based CL methods, which leverage the power of pretrained models, attribute forgetting either to (I) changes in parameters from the backbone when using the common prompt pool $\bm{P}$ for all tasks [49, 48] or to (II) the inherent mismatch between the models used at training and testing, as discussed and analyzed in [53, 40]. In this work, we complement these views by providing empirical studies to offer new insights about the reason for forgetting in this type of model, which arises from overlapping between old and new emerging class representations.
To uncover the new factor causing forgetting, we propose to consider the cases where the behavior of backbone w.r.t each learned task is nearly unchanged over time - meaning (I) and (II) roughly would not happen. Firstly, to eliminate concerns about changing learned parameters (I), we consider methods that propose using a distinct set of prompts $\bm{P}_{t}$ to a specific task $t$. Then, the remaining potential factor is the difference between the prompt chosen at inference time, which we refer as $\bm{P}_{\hat{t}}$, and the one used during training $\bm{P}_{t}$. Thus, we conduct experiments on HiDe-Prompt [44], which is the latest SOTA in prompt-based CL, to measure the differences between the features formed when using these two prompts. Particularly, we consider $\bm{x}\in\mathcal{D}_{1}$ that belongs to the first task and measure the L2 Wasserstein distance $W_{2}(Q,\hat{Q})$ [20] between $Q$ (i.e., the latent distribution corresponding to $\bm{P}_{t}$, which consists of $f_{\Phi,\bm{P}_{{t}}}(\bm{x})$ for $\bm{x}\sim\mathcal{D}_{1}$) and $\hat{Q}$ (i.e., the latent distribution corresponding $\bm{P}_{\hat{t}}$ after learning the last task, consists of $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$ for for $\bm{x}\sim\mathcal{D}_{1}$). The results in Table 1(a) show that the difference between these distributions is negligible when using Sup-21K as a pre-trained backbone, in stark contrast to the significant shift observed in DINO-1K.
For a detailed examination of Sup-21K, we propose an experimental setup to analyze the performance degradation of corresponding models over time. Specifically, we train a classifier $s(\bm{x})$ on the latent space $f_{\Phi,\bm{P}_{1}}(\bm{x})$ derived from the first task $\mathcal{D}_{1}$ to accurately classify the classes within this initial task. As the model undergoes continual training on a sequence of subsequent tasks, we assess the ability of $s(\bm{x})$ to classify instances from $\mathcal{D}_{1}$ based on the latent space representation $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$, defining this as within-task accuracy. This is compared to the true accuracy when using the classifier $h_{\psi}$. The results in Figure 1(b), show that within-task accuracy remains nearly stable, suggesting minimal shift between representations $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$ and $f_{\Phi,\bm{P}_{t}}(\bm{x})$. In contrast, Figure 1(c), illustrates a significant decrease in true accuracy, raising the question of whether we have overlooked additional factors contributing to final forgetfulness.
@@ 3.2 Forgetting in Prompt-based Methods
Additionally, considering the inference feature space of $f_{\Phi,\bm{P}_{\hat{t}}}(\bm{x})$, we can see that as more tasks arrive, the number of classes increase, making the space fuller and increasing the possibility of overlap between class distributions. To support this point, we illustrate the representation of some class images in Figure 2, as well as the corresponding t-SNE visualizations in Figure 4. In particular, after Task 1, we have representations of ”oak tree”, ”mouse” and ”porcupine” located in quite separate locations. However, when Task 2 and then Task 3 arrive, the appearance of ”willow tree” and ”pine tree” makes the latent space become fuller, and ”oak tree” no longer maintains the separation from the remaining classes as before, leading to a remarkable drop in model performance.
@@
Based on the evidence above, another key cause of catastrophic forgetting in prompt-based continual learning that should be recognized is the addition of new classes, which gradually fills the latent space and may overlap with existing ones. The overlap causes confusion in distinguishing between old and new classes, thereby reducing the performance of learned tasks over time. This is coincidently similar to the behavior of human memory, where although the old knowledge exists somewhere in the brain, still be confused due to the newly acquired information [1, 50, 26, 31, 51]. While existing CL methods emphasize the importance of representation learning to keep classes distinct [44, 24], none explicitly acknowledge the overlap between new and old tasks as a source of forgetting. Recognizing this motivates us to propose a novel method, focusing on identifying easily confused class pairs, thereby reducing forgetting and improving performance.
Figure 2: Problem: When new classes arrive, the latent space of a model for all tasks so far becomes fuller, and class representations tend to be overlapped, leading to performance degradation in old tasks. Our solution: In this work, we focus on separating easily confused classes whose concepts/labels lie in the same leaf group on a hierarchical taxonomy, suggested by expert knowledge (i.e., ChatGPT - Figure 3)
@@ 4 Proposed Method
In the previous section, we noted that increasing overlap in data representations as more tasks arrive is one of the main reasons leading to performance degradation (i.e., forgetting). Therefore, if we can identify easily confused classes/concepts, we will better enhance the distinguishability of models and thus reduce forgetting. Inspired by Cognitive Science studies, regarding the benefits of organizing information in a meaningful way (Section 4.1), we propose using expert knowledge to structure the concepts/labels of CL tasks in the corresponding hierarchical taxonomies. Interestingly, we find that concepts within the same group in each taxonomy tend to be visually and semantically similar, potentially causing more overlap in the latent space and confusion for the CL classifier. Motivated by this observation, we propose group-based contrastive learning in Section 4.2 to maximize the separability of these concepts, thereby mitigating forgetting over time. Additionally, to reinforce the effectiveness of this strategy, Section 4.3 presents an optimal transport-based strategy, which harnesses prior knowledge of pretrained backbones from a new perspective.